# Generalized Eckardt Points on del Pezzo surfaces of degree 1

### v1

@@ Abstract.
We study intersections of exceptional curves on del Pezzo surfaces of degree 1, motivated by questions in arithmetic geometry. Outside characteristics 2 and 3, at most 10 exceptional curves can intersect in a point. We classify the different ways in which 10 exceptional curves can intersect, construct a new families of surfaces with 10 exceptional curves intersecting in a point, and discuss strategies for finding more such examples.
@@ 1. Introduction
Del Pezzo surfaces are smooth projective algebraic surfaces with ample anticanonical divisor, and can be classified by their degree, an integer between 1 and 9. Over algebraically closed fields, these surfaces are rational (i.e., birationally equivalent to the projective plane). Over non-algebraically closed fields, this is not true in general, and the set $X(k)$ of rational points on a del Pezzo surface $X$ over such a field $k$ is far from fully understood. It is generally believed that del Pezzo surfaces that contain a rational point have ‘many’ rational points, and that these rational points are ‘well distributed’. There are several results that support this belief. Del Pezzo surfaces of degree at least 3 with a rational point are unirational (i.e., dominated by the projective plane) [Seg43, Seg51, Man74, Kol02, Pie12], for del Pezzo surfaces of degree 2 the same holds if they contain a rational point outside a closed subset [STVA14], and for del Pezzo surfaces of degree 1 unirationality is known when the surface contains a conic bundle [KM17]. If the field $k$ is infinite, then unirationality over $k$ implies that the set $X(k)$ is Zariski dense (i.e., not contained in a closed subset). There are several partial results on Zariski density of $X(k)$ for a del Pezzo surface of degree 1 [Ula08, Ula07, UT10, Jab12, SvL14, Des18, DW22]. Another measure of ‘many’ and ‘well distributed’ rational points for a variety $X$ over a number field $k$ is weak approximation (i.e., the set $X(k)$ of rational points is dense in the product of local points $\prod_{v\in\Omega_{k}}X(k_{v})$, where $\Omega_{k}$ is the set of places of $k$). Weak approximation is satisfied for all del Pezzo surfaces of degree at least 5 with a rational point, and shown not to hold in general for del Pezzo surfaces of lower degree. However, it is shown that del Pezzo surfaces of degree 3 and 4 with a rational point satisfy weak weak approximation (i.e., $X(k)$ is dense in $\prod_{v\in\Omega_{k}\setminus S}X(k_{v})$ for a finite set $S$) [SS91, SD01], and the same holds for del Pezzo surfaces of degree 2 with a general enough rational point [DSW23] and del Pezzo surfaces of degree 1 with a conic bundle under an extra condition [DS23]. As is evident from the known results, the arithmetic complexity of del Pezzo surfaces goes up as the degree goes down.
Over an algebraically closed field, a del Pezzo surface contains a fixed number of exceptional curves (which we often call lines), depending on the degree; the lower the degree, the more lines on the surface. These lines often come up when proving arithmetic results. For example, the results in [STVA14, DSW23] mentioned before require the existence of a rational point that is not contained in the intersection of 4 lines on a del Pezzo surface of degree 2. For del Pezzo surfaces of degree 1, the result in [DS23] requires that no 4 of the exceptional curves over the algebraic closure intersect in one point, the result in [SvL14] requires the existence of a point not contained in 6 or more lines, and in [DW22] one needs a point which is not torsion on its fiber of a corresponding elliptic surface, which implies that it can not be contained in 9 or more lines, as was shown in [DW].
@@
Following the notation established for del Pezzo surfaces of degrees 2 and 3, we call a point on a del Pezzo surface of degree 1 over a field with characteristic unequal to 2 and 3, that is contained in the intersection of 10 lines, a generalized Eckardt point. For characteristics 2 and 3 we define this to be the intersection point of 16 and 12 concurrent lines, respectively. The connection between points contained in many lines and the aforementioned arithmetic results prompts the following questions.
@@
In this paper we start an investigation towards answering Questions 1.1 and 1.2. In Section 2 we give the necessary background. In Section 3 we determine the different configurations in which 10 exceptional curves on a del Pezzo surface of degree 1 can pairwise intersect (Theorem 3.3 and Remark 3.4). Section 4 contains a discussion of how to determine which of these configurations can correspond to 10 exceptional curves that all intersect in one point, i.e., which configurations can correspond to generalized Eckardt points (Algorithm 4.6). In Section 5 we construct a family of surfaces that all contain a generalized Eckardt point (Theorem 5.1), and we study the same questions over finite fields in Section 6 (Theorem 6.5). We end with a short discussion on open questions in Section 7.
@@
Acknowledgements. This project originated during the workshop ‘Women in Numbers 6’, held in March 2023 at BIRS in Banff, Canada. We are grateful for the organisers and the organisation of BIRS. We thank Bianca Viray for useful discussions. Julie Desjardins was partially supported by an NSERC Discovery grant. Rosa Winter was supported by UKRI fellowship MR/T041609/1.
@@ 2.1. Del Pezzo surfaces of degree one and exceptional curves
A del Pezzo surface over a field $k$ is a smooth, projective, geometrically integral surface with ample anticanonical divisor. The degree of a del Pezzo surface is the self-intersection number of the canonical divisor, which is an integer between 1 and 9. Moreover, a del Pezzo surface over an algebraically closed field is either isomorphic to $\mathbb{P}^{1}\times\mathbb{P}^{1}$ when $d=8$ or to $\mathbb{P}^{2}$ blown up at $9-d$ points in general position otherwise ([Man74, Theorem 24.4]). Let $X$ be a del Pezzo surface of degree 1 over an algebraically closed field $k$, then $X$ is isomorphic to the blow up of $\mathbb{P}^{2}$ in eight points $P_{1},\ldots,P_{8}$ in general position. Let $\operatorname{Pic}(X)$ be the Picard group of $X$. Let $K_{X}$ be the class of the canonical divisor of $X$.
@@
$C\cdot C=-1,\text{ }C\cdot K_{X}=-1.$
The corresponding class $[C]\in\operatorname{Pic}(X)$ is called an exceptional class with
$[C]\cdot[C]=-1,\text{ }[C]\cdot K_{X}=-1.$
For $i\in\{1,\ldots,8\}$, let $[E_{i}]$ be the exceptional class of the exceptional curve above $P_{i}$, and $[L]$ be the class of the pullback of a line $\ell\subset\mathbb{P}^{2}$ not passing through any of the $P_{i}$. Once we realize $\operatorname{Pic}(X)$ as a lattice isomorphic to $\mathbb{Z}^{9}$, we can write down a basis as $\{[L],[E_{1}],\ldots,[E_{8}]\}$ ([Man74, Proposition 20.9.1]). Moreover, for $i,j\in\{1,\ldots,8\}$, $i\neq j$, we have
$[E_{i}]\cdot[E_{i}]=-1,\text{ }[E_{i}]\cdot[E_{j}]=0,\text{ }[L]\cdot[E_{i}]=0,\text{ }[L]\cdot[L]=1.$
@@
where $a$ is the degree of the corresponding curve in $\mathbb{P}^{2}$ and $b_{i}$ is the multiplicity of $P_{i}$. In this way, we obtain a vector $(a,b_{1},\ldots,b_{8})$ in $\mathbb{Z}^{9}$. Let $A$ be the set of 240 vectors obtained in this way, and let $C$ be the set of all exceptional classes in $\operatorname{Pic}(X)$. There is a one-to-one correspondence between $A$ and $C$ induced by the basis $L,E_{1},\ldots,E_{8}$ ([vLW23, Lemma 2.4, Remark 2.5]). Therefore we are able to study the intersection of exceptional curves (and classes) by carrying out computations in some vector spaces, either by magma or by hand.
@@ 2.2. The weighted graph on exceptional classes
An isomorphism between two weighted graphs $(V,D)$ and $\left(V^{\prime},D^{\prime}\right)$ with weight functions $\psi:D\longrightarrow$ $A$ and $\psi^{\prime}:D^{\prime}\longrightarrow A^{\prime}$, respectively, consists of a bijection $f$ between the sets $V$ and $V^{\prime}$ and a bijection $g$ between the sets $A$ and $A^{\prime}$, such that for any two vertices $v_{1},v_{2}\in V$, we have $\left\{v_{1},v_{2}\right\}\in D$ with weight $w$ if and only if $\{f\left(v_{1}\right),f\left(v_{2}\right)\}\in D^{\prime}$ with weight $g(w)$. We call the map $f$ an automorphism of $(V,D)$ if $(V,D)=\left(V^{\prime},D^{\prime}\right),\psi=\psi^{\prime}$, and $g$ is the identity on $A$.
@@
[…]
@@
We aim to study the intersection of exceptional curves in $X$, which is equivalent to studying the weighted graph $\Gamma$ on the $E_{8}$ root system. If two curves intersect, then the vertices corresponding to the exceptional curves are linked by edges of positive weights in $G$. The converse is not always true, however, as the exceptional curves in a clique can all pairwise intersect without intersecting in the same point! The possible positive weights that occur in $G$ are 1, 2, and 3, so the corresponding weights that occur in $\Gamma$ are 0, $-1$, and $-2$. The size of maximal cliques in $\Gamma$ gives an upper bound for the maximal possible number of exceptional curves in $X$ that intersect in one point.
@@
As mentioned in the introduction, the maximal number of concurrent exceptional curves on a del Pezzo surface of degree 3 is 3, and for a del Pezzo surface of degree 2, the number of exceptional curves that are concurrent in a point is at most 4. In both these cases, this number is the same as the maximal size of complete subgraphs of the intersection graph on the exceptional curves. But for del Pezzo surface of degree 1, things are more complicated. The maximal size of cliques in the graph $G$ is 16, hence this is also the maximal number of exceptional curves that are concurrent in a point on $X$ [vLW23, Proposition 2.8]. However, this is only sharp in characteristic 2; for char $k\neq 2$, the number of exceptional curves that are concurrent in a point is not given by $G$; it is at most ten outside characteristics 2 and 3, and at most 12 in characteristic 3 ([vLW23, Theorems 1.1, 1.2]). Both of these upper bounds are shown to be sharp except possibly in characteristic 5.
@@
    For $e_{1},e_{2}$ two exceptional classes in $\operatorname{Pic}(X)$ with $e_{1}\cdot e_{2}=3$, and $f$ a third exceptional class, we have $e_{1}\cdot f=1$ if and only if $e_{2}\cdot f=1$, and $e_{1}\cdot f=0$ if and only if $e_{2}\cdot f=2$.
The bi-anticanonical divisor ${-2K_{X}}$ induces a morphism
@@
such that $X$ can be realized as a double cover of a cone $Q$ in $\mathbb{P}^{3}$, ramified over a smooth sextic curve that does not contain the vertex of $Q$. By [CO99, Proposition 2.6], $\varphi$ maps an exceptional curve on $X$ bijectively to a smooth conic which is the intersection of $Q$ with a hyperplane in $\mathbb{P}^{3}$ not containing the vertex of $Q$. Conversely, for a hyperplane section $H$ of $Q$ not containing the vertex of $Q$ and tri-tangent to the smooth sextic curve, the pullback $\varphi^{\ast}H=e_{1}+e_{2}$ under $\varphi$ is the sum of two exceptional curves intersecting with multiplicity 3. We call $e_{1}$ the partner of $e_{2}$. Thus every point on $H$ has two preimages, except for the points with preimages in $e_{1}\cap e_{2}$.

### v2

@@ Abstract.
We study intersections of exceptional curves on del Pezzo surfaces of degree 1, motivated by questions in arithmetic geometry. Outside characteristics 2 and 3, at most 10 exceptional curves can intersect in a point. We classify the different ways in which 10 exceptional curves can intersect, construct a new family of surfaces with 10 exceptional curves intersecting in a point, and discuss strategies for finding more such examples.
@@ 1. Introduction
Del Pezzo surfaces are smooth projective algebraic surfaces with ample anticanonical divisor, and can be classified by their degree, which is defined as the self-intersection of the canonical divisor; an integer between 1 and 9. Both the geometric as well as the arithmetic complexity of del Pezzo surfaces goes up as the degree goes down (see also the Motivation from the rational points perspective below). Over an algebraically closed field, a del Pezzo surface contains a fixed number of exceptional curves (which we often call lines), depending on the degree; the lower the degree, the more lines on the surface.
@@
Following the terminology established for del Pezzo surfaces of degrees 2 and 3, we call a point on a del Pezzo surface of degree 1 over a field with characteristic unequal to 2 and 3, that is contained in the intersection of 10 lines, a generalized Eckardt point. For characteristics 2 and 3 we define this to be the intersection point of 16 and 12 lines, respectively. The knowledge we have about del Pezzo surfaces of degree at least 2 prompts the following questions.
@@
In this paper we start an investigation towards answering Questions 1.1 and 1.2. Our main contribution is the following result (c.f. Theorem 3.3):
Theorem 1.3.
Outside characteristic 2, there are 9 possible configurations for 10 exceptional curves on a del Pezzo surface of degree 1 to intersect in one point, and they are listed in Figures 2 and 3. In characteristic 0, at most 7 of the configurations (1, 2, 3, 5, 6, 8 in Figure 2 and the configuration in Figure 3) can correspond to a generalized Eckardt point on a del Pezzo surface of degree 1.
Theorem 3.3 is a first step in answering Questions 1.1 and 1.2, as it tells us which configurations of intersecting lines to count. As mentioned, only two examples of (families of) del Pezzo surfaces of degree 1 with a generalized Eckardt points can be found in the current literature. In [SvL14, Example 4.1], the authors construct a family of del Pezzo surfaces of degree 1 that all contain 10 exceptional curves that intersect in one point, whose intersection graph is isomorphic to number 5 in Figure 2. In [vLW23, Example 5.2] the authors construct a del Pezzo surface of degree 1 over a field of characteristic unequal to 2 with 10 exceptional curves that intersect in one point, whose intersection graph is isomorphic to the one in Figure 3. We extend this to a whole family of such examples in Theorem 5.1.
The strategy used in order to obtain the examples of Theorem 5.1 is explained in detail with Algorithm 4.6. This strategy combined with combinatorial computations proves that there are no del Pezzo surface of degree 1 with a generalized Eckardt point corresponding to the configurations in Figure 2 over the finite field $\mathbb{F}_{p}$ with $p\in\{1,\ldots,17,23\}$ (Theorems 6.5 and 6.6) and allows us to find an example of a del Pezzo surface of degree 1 with a generalized Eckardt point corresponding to configuration 8 over $\mathbb{F}_{19}$ (Theorem 6.7). We are quite confident that our strategy will lead to more results of the kind we are presenting in Sections 5 and 6.
Motivation from the rational points perspective. One motivation to study Questions 1.1 and 1.2 comes from the arithmetic of del Pezzo surfaces. Over algebraically closed fields, these surfaces are rational (i.e., birationally equivalent to the projective plane). Over non-algebraically closed fields, this is not true in general, and the set $X(k)$ of rational points on a del Pezzo surface $X$ over such a field $k$ is far from fully understood. It is generally believed that del Pezzo surfaces that contain a rational point have ‘many’ rational points, and that these rational points are ‘well distributed’. There are several results that support this belief. Del Pezzo surfaces of degree at least 3 with a rational point are unirational (i.e., dominated by the projective plane) [Seg43, Seg51, Man74, Kol02, Pie12], for del Pezzo surfaces of degree 2 the same holds if they contain a rational point outside a closed subset [STVA14], and for del Pezzo surfaces of degree 1 unirationality is known when the surface contains a conic bundle [KM17]. If the field $k$ is infinite, then unirationality over $k$ implies that the set $X(k)$ is Zariski dense (i.e., not contained in a closed subset). There are several partial results on Zariski density of $X(k)$ for a del Pezzo surface of degree 1 [Ula08, Ula07, UT10, Jab12, SvL14, Des18, DW22]. Another measure of ‘many’ and ‘well distributed’ rational points for a variety $X$ over a number field $k$ is weak approximation (i.e., the set $X(k)$ of rational points is dense in the product of local points $\prod_{v\in\Omega_{k}}X(k_{v})$, where $\Omega_{k}$ is the set of places of $k$). Weak approximation is satisfied for all del Pezzo surfaces of degree at least 5 with a rational point, and shown not to hold in general for del Pezzo surfaces of lower degree. However, it is shown that del Pezzo surfaces of degree 3 and 4 with a rational point satisfy weak weak approximation (i.e., $X(k)$ is dense in $\prod_{v\in\Omega_{k}\setminus S}X(k_{v})$ for a finite set $S$) [SS91, SD01], and the same holds for del Pezzo surfaces of degree 2 with a general enough rational point [DSW23] and del Pezzo surfaces of degree 1 with a conic bundle under an extra condition [DS23].
The configuration of lines often come up when proving these arithmetic results. For example, the results in [STVA14, DSW23] require the existence of a rational point that is not contained in the intersection of 4 lines on a del Pezzo surface of degree 2. For del Pezzo surfaces of degree 1, the result in [DS23] requires that no 4 of the exceptional curves over the algebraic closure intersect in one point, the result in [SvL14] requires the existence of a point not contained in 6 or more lines, and in [DW22] one needs a point which is not torsion on its fiber of a corresponding elliptic surface, which implies that it can not be contained in 9 or more lines, as was shown in [DW].
Outline of the paper
@@
In Section 2 we give the necessary background. In Section 3 we determine the different configurations in which 10 exceptional curves on a del Pezzo surface of degree 1 can pairwise intersect (Theorem 3.3 and Remark 3.4). Section 4 contains a discussion of how to determine which of these configurations can correspond to 10 exceptional curves that all intersect in one point, i.e., which configurations can correspond to generalized Eckardt points (Algorithm 4.6). In Section 5 we construct a family of surfaces that all contain a generalized Eckardt point (Theorem 5.1), and we study the same questions over finite fields in Section 6 (Theorems 6.5, 6.6, 6.7). We end with a short discussion on open questions in Section 7.
@@
Acknowledgements. This project originated during the workshop ‘Women in Numbers 6’, held in March 2023 at BIRS in Banff, Canada. We are grateful for the organisers and the organisation of BIRS. We thank Bianca Viray for useful discussions. We are grateful to the anonymous referee for useful comments that improved the quality of the paper. Julie Desjardins was partially supported by an NSERC Discovery grant. Rosa Winter was supported by UKRI fellowship MR/T041609/2.
@@ 2.1. Del Pezzo surfaces of degree one and exceptional curves
A del Pezzo surface over a field $k$ is a smooth, projective, geometrically integral surface with ample anticanonical divisor. The degree of a del Pezzo surface is the self-intersection number of the canonical divisor, which is an integer between 1 and 9. Moreover, a del Pezzo surface over an algebraically closed field is either isomorphic to $\mathbb{P}^{1}\times\mathbb{P}^{1}$ when $d=8$ or to $\mathbb{P}^{2}$ blown up at 9d9-d points in general position otherwise [Man74, Theorem 24.4]. Let $X$ be a del Pezzo surface of degree 1 over an algebraically closed field $k$, then $X$ is isomorphic to the blow up of $\mathbb{P}^{2}$ in eight points $P_{1},\ldots,P_{8}$ in general position. Let $\operatorname{Pic}(X)$ be the Picard group of $X$. Let $K_{X}$ be a fixed divisor in the canonical class of $X$.
@@
CC=1, $\cdot C=-1,\text{ }C\cdot K_{X}$=-1.
The corresponding class $C]\in\operatorname{Pic}(X)$ is called an exceptional class with
[C][C]=1, $\cdot[C]=-1,\text{ }[C]\cdot[K_{X}]$=-1.
For $i\in\{1,\ldots,8\}$, let $E_{i}]$ be the exceptional class of the exceptional curve above $P_{i}$, and $[L]$ be the class of the pullback of a line $\ell\subset\mathbb{P}^{2}$ not passing through any of the $P_{i}$. Once we realize $\operatorname{Pic}(X)$ as a lattice isomorphic to $\mathbb{Z}^{9}$, we can write down a basis as $\{[L],[E_{1}],\ldots,[E_{8}]\}$ [Man74, Proposition 20.9.1]. Moreover, for $\in\{1,\ldots,8\}, iji\neq j$, we have
$E_{i}]\cdot[E_{i}]=-1,\text{ }[E_{i}]\cdot[E_{j}]=0,\text{ }[L]\cdot[E_{i}]=0,\text{ }[L]\cdot[L]=1.$
@@
where $a$ is the degree of the corresponding curve in $\mathbb{P}^{2}$ and $b_{i}$ is the multiplicity of $P_{i}$. In this way, we obtain a vector $(a,b_{1},\ldots,b_{8})$ in $\mathbb{Z}^{9}$. Let $A$ be the set of 240 vectors obtained in this way, and let $C$ be the set of all exceptional classes in $\operatorname{Pic}(X)$. There is a one-to-one correspondence between $A$ and $C$ induced by the basis $L,E_{1},\ldots,E_{8}$ [vLW23, Lemma 2.4, Remark 2.5]. Therefore we are able to study the intersection of exceptional curves (and classes) by carrying out computations in some vector spaces, either by magma or by hand.
@@ 2.2. The weighted graph on exceptional classes
An isomorphism between two weighted graphs $(V,D)$ and $\left(V^{\prime},D^{\prime}\right)$ with weight functions $\psi:D\longrightarrow$ AA and $\psi^{\prime}:D^{\prime}\longrightarrow A^{\prime}$, respectively, consists of a bijection $f$ between the sets $V$ and $V^{\prime}$ and a bijection $g$ between the sets $A$ and $A^{\prime}$, such that for any two vertices $v_{1},v_{2}\in V$, we have $\left\{v_{1},v_{2}\right\}\in D$ with weight $w$ if and only if $\{f\left(v_{1}\right),f\left(v_{2}\right)\}\in D^{\prime}$ with weight $g(w)$. We call the map $f$ an automorphism of $(V,D)$ if $(V,D)=\left(V^{\prime},D^{\prime}\right),\psi=\psi^{\prime}$, and $g$ is the identity on $A$.
@@
Let $X$ be a del Pezzo surface of degree 1 with canonical divisor $K_{X}$, and let $K_{X}^{\perp}$ be the orthogonal complement of $K_{X}$ in $\operatorname{Pic}(X)$. Let $\langle\cdot,\cdot\rangle$ be the negative of the intersection pairing on $\operatorname{Pic}(X)$ and note that $\langle\cdot,\cdot\rangle$ induces an inner product on $\mathbb{R}\otimes_{\mathbb{Z}}\operatorname{Pic}(X)$ which realizes $K_{X}^{\perp}$ as Euclidean space. Vectors in $K_{X}^{\perp}$ with length $\sqrt{2}$, i.e., classes in $K_{X}^{\perp}$ with self-intersection 2-2, form a root system of type $E_{8}$ [Man74, Theorem 23.9]:
@@
We aim to study the intersection of exceptional curves in $X$. If two such curves intersect, then the corresponding vertices in $G$ are connected by edges of positive weights. The possible positive weights that occur in $G$ are 1, 2, and 3, so the corresponding weights that occur in $\Gamma$ are 0, 1-1, and 2-2. Therefore, the size of maximal cliques in $\Gamma$ with edges of these weights gives an upper bound for the maximal possible number of lines on $X$ that go through one point. Note that this upper bound does not need to be sharp.
@@
As mentioned in the introduction, the maximal number of concurrent exceptional curves on a del Pezzo surface of degree 3 is 3, and for a del Pezzo surface of degree 2, the number of exceptional curves that are concurrent in a point is at most 4. In both these cases, this number is the same as the maximal size of complete subgraphs of the intersection graph on the exceptional curves. But for del Pezzo surface of degree 1, things are more complicated. The maximal size of cliques in the graph $G$ is 16, hence this is also the maximal number of exceptional curves that are concurrent in a point on $X$ [vLW23, Proposition 2.8]. However, this is only sharp in characteristic 2; for char $k\neq 2$, the number of exceptional curves that are concurrent in a point is not given by the maximal size of cliques in $G$; it is at most 10 outside characteristics 2 and 3, and at most 12 in characteristic 3 [vLW23, Theorems 1.1, 1.2]. Both of these upper bounds are shown to be sharp except possibly in characteristic 5.
@@
    For $e_{1},e_{2}$ two exceptional classes in $\operatorname{Pic}(X)$ with $e_{1}\cdot e_{2}=3$, and $f$ a third exceptional class, we have $e_{1}\cdot f+e_{2}\cdot f=2$.
The bi-anticanonical divisor $2K_{X}$} induces a morphism
@@
such that $X$ can be realized as a double cover of a cone $Q$ in $\mathbb{P}^{3}$, ramified over a sextic curve [CO99, Proposition 2.3]. Moreover, by [CO99, Proposition 2.6], $\varphi$ maps an exceptional curve on $X$ bijectively to a smooth conic which is the intersection of $Q$ with a hyperplane in $\mathbb{P}^{3}$ not containing the vertex of $Q$. Conversely, for a hyperplane section $H$ of $Q$ not containing the vertex of $Q$ and tri-tangent to the smooth sextic curve, the pullback $\varphi^{\ast}H=e_{1}+e_{2}$ under $\varphi$ is the sum of two exceptional curves intersecting with multiplicity 3. We call $e_{1}$ the partner of $e_{2}$. Thus every point on $H$ has two preimages, except for the points with preimages in $e_{1}\cap e_{2}$.